Given the equation 3x-12y=24 Find the x-axis intercept ( , ) and the y-axis intercept ( , ) of the line through the points.
Answer by saltzy57(8)

Y intercept form....Y=mX+b
Now rewrite your equation in Y intercept form
12y=3x-24....solve for y...
y=3/12(x)-2 simplify
y=1/4*x -2
for Y axis intercept, let x=0...Y=-2
for X axis intercept ley Y=0... X=8